# LD-SDM: Language-Driven Hierarchical Species Distribution Modeling

### v1

@@ 4.2 Proximity-Aware Evaluation
Loss Env. Method MAP(pa) $\uparrow$ $d_{\text{PWCD}}$downarrow MAP(pa) $\uparrow$ $d_{\text{PWCD}}$downarrow MAP(pa) $\uparrow$ $d_{\text{PWCD}}$downarrow (Mean $R^{2}$)
$\mathcal{L}_{\text{AN-full}}$ [15] SINR [15] 62.64 0.177 60.28 0.178 - - 0.446
$\mathcal{L}_{\text{ME-full}}$ [49] SINR 61.88 0.189 58.81 0.187 - - 0.518
$\mathcal{L}_{\text{ASL}}$ [32] SINR 67.11 0.152 64.55 0.155 - - 0.619
$\mathcal{L}_{\text{RAL}}$ [31] SINR 62.02 0.291 60.42 0.297 - - 0.536
$\mathcal{L}_{\text{AN-full}}$ SIREN(SH) [34] 61.99 0.203 60.67 0.215 - - 0.562
@@
Text Encoder encoding dim MAP(pa) $\uparrow$ $d_{PWCD}$downarrow MAP(pa) $\uparrow$ $d_{PWCD}$downarrow MAP(pa) $\uparrow$ $d_{PWCD}$downarrow (Mean $R^{2}$) TFLOPS
@@
4.3 Training and Evaluation
We focus on bird species occurrence data which is abundantly available yet challenging to model as compared to other species due to complex migration patterns. We compiled 131.7 million observations of 4141 species of birds using the GBIF (gbif.org) platform which includes data from various community science platforms. This data is collected for the years 2012-2020. For evaluation, we use curated research-grade observations from the iNaturalist [44] and eBird [37] for 2021 and 2022. These contain 17.9 million and 17.2 million observations respectively. Using the same platform, we created a dataset of 84 species that are unseen (not present in training data) and rare ($<$ 1000 observations) for zero-shot performance evaluation. This contains a total of 26k observations. These platforms are useful since errors in misclassification are highly mitigated due to community consensus [24], hence being a strong representative of the ground truth. Further, we use the same procedure as described in [15] for the geo-feature regression task.
Figure 3: Predicted Range Maps (Probability Values). The predictions are generated at a spatial resolution of 0.2 degrees. The best performing models are used in each case. For the SIREN(SH) and SINR models, predictions are generated at the center of each raster cell. The visualizations show that our model performs better than other models at resolving higher frequencies over the sphere and localizing the species distribution.
@@ 4.4 Results
Quantitative Evaluation. In Table 1, we report the performance of all the models considered in this work. LD-SDM has shown better performance than the other models on the task of species range prediction. The zero-shot performance of LD-SDM on unseen species is reasonably good. Although it has an increased false-positive rate, it can be attributed to the model using higher-level taxonomic relationships. $\mathcal{L}_{\text{RAL}}$ tends to produce overconfident predictions as noticeable from the high MAP(pa) scores corresponding to high $d_{PWCD}$ scores (especially for SINR and SIREN(SH)). On average, $\mathcal{L}_{\text{ASL}}$ loss is likely to be better performing than the other losses over all the models. For SINR, $\mathcal{L}_{\text{ASL}}$ performs the best as compared to the other losses by a large margin. This is also evident in the geo-feature regression task. Our model has also produced competitive scores on the task of geo-feature regression. The coordinate-only models perform nearly as well as models incorporating environmental features. So, dropping environmental features could result in faster training times.
LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
Taxonomic Level MAP(pa) $\uparrow$ $d_{PWCD}$downarrow MAP(pa) $\uparrow$ $d_{PWCD}$downarrow MAP(pa) $\uparrow$ $d_{PWCD}$downarrow
Class 97.10 0.117 95.2 0.204 97.52 0.163
Order 56.15 0.450 55.94 0.472 58.31 0.414
Family 38.18 0.659 36.47 0.688 40.13 0.632
Genus 24.95 0.710 24.97 0.711 27.86 0.692
@@
Backbone M MAP(pa) $\uparrow$ $d_{PWCD}$downarrow MAP(pa) $\uparrow$ $d_{PWCD}$downarrow
@@
GBIF’21 GBIF’22
encoder dim MAP(pa) $\uparrow$ $d_{PWCD}$downarrow MAP(pa) $\uparrow$ $d_{PWCD}$downarrow
64 68.55 0.147 67.28 0.149
128 73.88 0.174 71.45 0.182
256 73.67 0.178 71.82 0.188
(a) SFNO width
GBIF’21 GBIF’22
encoder layers MAP(pa) $\uparrow$ PWCD $\downarrow$ MAP(pa) $\uparrow$ PWCD $\downarrow$
2 73.88 0.174 71.45 0.182
3 73.39 0.171 72.11 0.174
4 73.67 0.174 72.00 0.180
(b) SFNO depth
Table 5: SFNO ablation in LD-SDM. All experiments run with the $\mathcal{L}_{\text{ASL}}$ loss and the LLaMA-2-70B variant.
@@
Qualitative Examples. In Figure 3, we show global-scale species distribution maps generated for five bird species. To make fair comparisons, we used the best-performing coordinate-only models to ensure no bias is added into models from the environmental features. Although SINR and SIREN(SH) could generate species distribution maps at any spatial resolution, they tend to learn a very low-frequency function over the sphere as evident from the blob-like maps. SIREN(SH) produces a lot of false positive predictions possibly due to the limitation of the number of Legendre polynomials it can fit. Our model on the other hand learns to predict species ranges at a more fine-grained level. This is evident from the visualization of the geographic location features learned by the location encoder of each model (shown in Appendix D). Fine-grained geographic features are important for better geolocalization of species. SINR and SIREN(SH) tend to learn similar features over large geographic regions while LD-SDM produces higher variability in the features at a finer scale.
@@
Effect of LLaMA-2 variants on LD-SDM. We assess the impact of using different variants of LLaMA-2 for LD-SDM in Table 2. In particular, we compare the 7B, 13B, and 70B variants. Surprisingly, the 13B variant performs the worst. This is possibly due to the differences in the training data of the variants. Empirically, this can be explained by examining the T-SNE plots of text embeddings. In Appendix A, we present T-SNE plots along with mean intra-cluster and inter-cluster distances of text embeddings. Ideally, both intra-cluster and inter-cluster distances should be high to ensure separability within a particular taxonomic class and between taxonomic classes respectively. 13B variant has the poorest separability at every taxonomic rank. The 7B and 70B variants have almost identical separatabilities. The impressive performance of the 70B variant can be attributed to its high dimensionality of the species-specific encoding. Although 70B has noticeably higher computational overheads, it is compensated by precomputing the text embeddings before training. Overall, all the variants are able to perform better than the state-of-the-art.
Taxonomic range mapping using LLaMA-2 variants. We perform range prediction over the taxonomic ranks: class, order, family, and genus. The procedure is the same as for generating species range maps. We precompute the embeddings for each rank as shown in Figure 1. The range map for the rank class represents the unconditional distribution of finding any bird species in the world. As one moves down in the hierarchy, the range maps become more and more fine-grained. As expected, the performance of LD-SDM drops as one moves down in the hierarchy (Table  3). This can be explained by examining the increasing $d_{\text{PWCD}}$ scores. The T-SNE plots of the embeddings of various ranks in Appendix A help in better understanding the increase in false positive rates as one moves down in the hierarchy. As it can be seen, the intra-cluster distances decrease substantially when moving down in the hierarchy. As a result, it becomes challenging for the model to distinguish between birds for example of genera belonging to the same family. On the other hand, when training in a fully supervised manner at the species rank, the model is able to achieve impressive performance. This shows the importance of supervision in SDM.
@@
Spherical Harmonics are important for SDM. We compare the performance of LD-SDM when using different backbones in the location encoder network (Table 4). All the models are trained from scratch using the LLaMA-2-70B variant. We use the same training inputs including the environmental features. All the models are trained using the $\mathcal{L}_{\text{ASL}}$ loss. Notice that SFNO gives the best performance over Segformer and DINOv2. With SFNO, we observe a significant gain in performance and a reduction in false positives. Overall, a robust geographic feature extractor is beneficial for SDM and SFNO proves to be one such architecture.
Lightweight SFNO performs surprisingly well. In Table 5, we show the effect on the performance of species range prediction due to different components of the SFNO. It is noteworthy that an encoder dimension of 128 for SFNO yields results on par with an encoder dimension of 256 while reducing computational complexity. On the other hand, having fewer encoder layers in SFNO is beneficial since performance gain from adding additional encoder layers is marginal. In conclusion, a lightweight SFNO is a strong geographic feature extractor that can be used in various downstream global-scale geospatial tasks.
@@ 5 Conclusions
Appendix A Species-Specific Embeddings
@@
Appendix B Probability Weighted Chamfer Distance
@@ C.1 Data
To avoid bias, we excluded species with less than 100 observations. The evaluation datasets only contain observations from iNaturalist and eBird. For zero-shot evaluation, we selected the species that were not present in the training dataset. Further, we filtered out species with more than 1000 observations to focus on rare species. Table 7 shows the details of each dataset.
Table 7: Number of observations in each dataset compiled in this work.
Dataset Year #Observations #species
Train 2012-2020 131.7M 4141
Test 2021 17.9M 3960
Test 2022 17.2M 4010
Zero-Shot 2021-2022 26k 84
@@
C.2 Models
We experimented with three different SDMs: 1) SINR [15]; 2) SIREN(SH) [34]; LD-SDM (ours). For the baseline models, we used the same hyperparameters as reported by authors. Tables 89 and 10, show the hyperparameters used for each model. We use $L=20$ for SIREN(SH) since experiments with larger $L$ were computationally infeasible.
C.3 Losses
@@
Appendix D Location Embeddings
@@
Appendix E Limitations
As pointed out in [15], the data obtained from the process of crowdsourcing is biased towards regions with heavy human traffic. Most often, rich biodiverse regions are underrepresented limiting the performance of SDMs. In this work, we do not explicitly deal with the observation bias in presence-only SDM. However, our goal was to incorporate taxonomic information into SDM to enhance unseen species discovery.
Previous studies have evaluated species SDMs using expert range maps, which are only available for a limited number of species. Our proposed evaluation metric allows for the evaluation of SDMs using crowdsourced data, which provides a more comprehensive representation of numerous species. However, we note that none of these methods are completely free of errors. Yet, with our proposed evaluation metric, innovative methods could be developed for semi-automatic evaluation of SDMs in the future, to strike a balance between accuracy and efficiency.

### v2

@@ 4.3 Training and Evaluation
Loss Env. Method MAP(pa) $\uparrow$ $d_{\text{PWCD}}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{\text{PWCD}}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{\text{PWCD}}$ $\downarrow$ (Mean $R^{2}$)
$\mathcal{L}_{\text{AN-full}}$ SINR 62.64 0.177 60.28 0.178 - - 0.446
$\mathcal{L}_{\text{ME-full}}$ SINR 61.88 0.189 58.81 0.187 - - 0.518
$\mathcal{L}_{\text{ASL}}$ SINR 67.11 0.152 64.55 0.155 - - 0.619
$\mathcal{L}_{\text{RAL}}$ SINR 62.02 0.291 60.42 0.297 - - 0.536
$\mathcal{L}_{\text{AN-full}}$ SIREN(SH) 61.99 0.203 60.67 0.215 - - 0.562
@@
Text Encoder encoding dim MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ (Mean $R^{2}$) TFLOPS
@@
Figure 4: Predicted Range Maps (Probability Values). The predictions are generated at a spatial resolution of 0.2 degrees. For the SIREN(SH) and SINR models, predictions are generated at the center of each raster cell. The visualizations show that our model performs better than other models at resolving higher frequencies over the sphere and localizing the species distribution.
@@ 4.4 Results
Quantitative Evaluation. In Table 2, we report the performance of all the models considered in this work. LD-SDM has shown better performance than the other models on the task of species range prediction. The zero-shot performance of LD-SDM on unseen species is reasonably good. Although it has an increased false-positive rate, it can be attributed to the model using higher-level taxonomic relationships. Our metric is sensitive to model output confidence ranges. SINR and SIREN(SH) output likelihood values in 0.0-0.3 range, while our model produces confidence in 0.0-1.0 range. This results in lower false positive and true positive rates for the former models than ours. $\mathcal{L}_{\text{RAL}}$ tends to produce overconfident predictions as noticeable from the high MAP(pa) scores corresponding to high $d_{PWCD}$ scores. On average, $\mathcal{L}_{\text{ASL}}$ loss is likely to be better performing than the other losses over all the models. For SINR, $\mathcal{L}_{\text{ASL}}$ performs the best as compared to the other losses by a large margin. This is also evident in the geo-feature regression task. Our model has also produced competitive scores on the task of geo-feature regression. The coordinate-only models perform nearly as well as models incorporating environmental features. So, dropping environmental features could result in faster training.
@@
Backbone M MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$
@@
Qualitative Examples. In Figure 4, we show global-scale species distribution maps generated for five bird species. To make fair comparisons, we used the best-performing coordinate-only models to ensure no bias is added to models from the environmental features. Although SINR and SIREN(SH) could generate species distribution maps at any spatial resolution, they tend to learn a very low-frequency function over the sphere as evident from the blob-like maps. SIREN(SH) produces a lot of false positive predictions possibly due to the limitation of the number of Legendre polynomials it can fit. Our model on the other hand learns to predict species ranges at a more fine-grained level. This is evident from the visualization of the geographic location features learned by the location encoder of each model (shown in Appendix D). Fine-grained geographic features are important for better geolocalization of species. SINR and SIREN(SH) tend to learn similar features over large geographic regions while LD-SDM produces higher variability in the features at a finer scale. The learned features could be used for other ecological mapping tasks that require geography-aware features.
@@
Effect of LLaMA-2 variants on LD-SDM. We assess the impact of using different variants of LLaMA-2 for LD-SDM in Table 3. In particular, we compare the 7B, 13B, and 70B variants. Surprisingly, the 13B variant performs the worst. This is possibly due to the differences in the training data of the variants. Empirically, this can be explained by examining the T-SNE plots of text embeddings. In Appendix A, we present T-SNE plots along with mean intra-cluster and inter-cluster distances of text embeddings. Ideally, both intra-cluster and inter-cluster distances should be high to ensure separability within a particular taxonomic class and between taxonomic classes respectively. 13B variant has the poorest separability at every taxonomic rank. The 7B and 70B variants have almost identical separatabilities. The impressive performance of the 70B variant can be attributed to its high dimensionality of the species-specific encoding. Although 70B has noticeably higher computational overheads, it is compensated by precomputing the text embeddings before training. Overall, all the variants can perform better than the state-of-the-art.
@@
Spherical Harmonics are important for SDM. We compare the performance of LD-SDM when using different backbones in the location encoder network (Table 4). All the models are trained from scratch using the LLaMA-2-70B variant. We use the same training inputs including the environmental features. This is done since positional encodings are implicitly incorporated by the Segformer and DINOv2 architectures. All the models are trained using the $\mathcal{L}_{\text{ASL}}$ loss. Notice that SFNO gives the best performance over Segformer and DINOv2. With SFNO, we observe a significant gain in performance and a reduction in false positives. Overall, a robust geographic feature extractor is beneficial for SDM and SFNO proves to be one such architecture.
@@ 5 Conclusions
6 Species-Specific Embeddings
@@
7 Probability Weighted Chamfer Distance
@@ 8.1 Data
To avoid bias, we excluded species with less than 100 observations. The evaluation datasets only contain observations from iNaturalist and eBird. For zero-shot evaluation, we selected the species that were not present in the training dataset. Further, we filtered out species with more than 1000 observations to focus on rare species.
@@
8.2 Models
We experimented with three different SDMs: 1) SINR [14]; 2) SIREN(SH) [37]; LD-SDM (ours). For the baseline models, we used the same hyperparameters as reported by authors. Tables 67 and 8, show the hyperparameters used for each model. We use $L=20$ for SIREN(SH) since experiments with larger $L$ were computationally infeasible.
8.3 Losses
@@
Table 9: SFNO ablation in LD-SDM. We experiment on various values of hidden encoder dimension and the number of encoder layers in SFNO. All experiments run with the $\mathcal{L}_{\text{ASL}}$ loss and the LLaMA-2-70B variant.
GBIF’21 GBIF’22
Encoder Dim. MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$
64 68.55 0.147 67.28 0.149
128 73.88 0.174 71.45 0.182
256 73.67 0.178 71.82 0.188
Encoder Layers MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$
2 73.88 0.174 71.45 0.182
3 73.39 0.171 72.11 0.174
4 73.67 0.174 72.00 0.180
8.4 SFNO
In Table 9, we show the effect on the performance of species range prediction due to different components of the SFNO. It is noteworthy that an encoder dimension of 128 for SFNO yields results on par with an encoder dimension of 256 while reducing computational complexity. On the other hand, having fewer encoder layers in SFNO is beneficial since performance gain from adding additional encoder layers is marginal. In conclusion, a lightweight SFNO is a strong geographic feature extractor that can be used in various downstream global-scale geospatial tasks.
LLaMA-2-7B LLaMA-2-13B LLaMA-2-70B
Taxonomic Level MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$ MAP(pa) $\uparrow$ $d_{PWCD}$ $\downarrow$
Class 97.10 0.117 95.2 0.204 97.52 0.163
Order 56.15 0.450 55.94 0.472 58.31 0.414
Family 38.18 0.659 36.47 0.688 40.13 0.632
Genus 24.95 0.710 24.97 0.711 27.86 0.692
8.5 LLaMA-2 Variants
We perform range prediction over the taxonomic ranks: class, order, family, and genus. The procedure is the same as for generating species range maps. We precompute the embeddings for each rank as shown in Figure 1. The range map for the rank class represents the unconditional distribution of finding any bird species in the world. As one moves down in the hierarchy, the range maps become more and more fine-grained. As expected, the performance of LD-SDM drops as one moves down in the hierarchy (Table  10). This can be explained by examining the increasing $d_{\text{PWCD}}$ scores. The T-SNE plots of the embeddings of various ranks in Appendix A help in better understanding the increase in false positive rates as one moves down in the hierarchy. As it can be seen, the intra-cluster distances decrease substantially when moving down in the hierarchy. As a result, it becomes challenging for the model to distinguish between birds for example of genera belonging to the same family.
9 Location Embeddings